# Internal and String Stability of an Observer-based Controller for Vehicle Platooning under the MPF Topology

### v2

@@ Abstract
In this paper, we study the internal stability and string stability of a vehicle platoon under the constant time headway spacing (CTHS) policy and the multiple-predecessor-following (MPF) vehicle-to-vehicle information flow topology. More specifically, we depart from the conventional Proportional-Integral-Derivative (PID) controller design for such systems and we propose the design of an observer-based controller. For designing our observer-based controller, we first design a distributed observer, with which each follower estimates their position, speed and acceleration error with respect to the leader. The observer is designed by means of constructing an observer matrix whose parameters should be determined. Next, we simplify the design of the matrix of the observer in such a way that the design boils down to choosing a single scalar value; this design further simplifies the structure of the controller, whose simplicity enables the derivation of string stability conditions by means of a frequency response method. Subsequently, the string stability conditions for a given time headway, are transformed to conditions for the controller parameters. We obtain controller parameters that satisfy the stability conditions by designing a novel heuristic search algorithm. Furthermore, we extend the search algorithm by incorporating a bisection-like algorithm, which allows to obtain (within some deviation tolerance) the minimum available value of the time headway. Finally, we provide insights about how to finalize the observer-based controller parameters from above algorithms to avoid the peaking phenomenon. The performance of the proposed observer-based controller, which guarantees internal and string stability, is demonstrated via illustrative examples. Additionally, a comparison with a widely-used PID controller for MPF topology shows that our proposed observer-based controller has better convergence performance.
@@
Vehicle platooning, varying-speed leader, observer, string stability, V2V communication.
@@ I Introduction
    In order to improve platooning control performance under the MPF topology, a distributed observer-based controller is proposed, which can also ensure the internal and string stability of the platoon system One simulated example in Sec. VI-A demonstrates that this distributed observer-based controller achieves a better platooning control performance compared to the distributed PID controller of [6]. The key point in analyzing string stability is to propose a new calculation mechanism for deriving string stability conditions.
@@
    A bisection-like algorithm is proposed to obtain the minimum (within a tolerance) acceptable value of the time headway that can guarantee string stability. Another simulated example in Sec. VI-C demonstrates that the obtained minimum time headway is smaller than the minimum one from [6] under the same platooning model conditions.
The rest of this paper is organized as follows. Sec. II gives some notations and mathematical preliminaries. Sec. III presents the vehicle model, the spacing policy and the control objective. In Sec. IV, a distributed observer-based controller is proposed with the internal stability analysis. String stability of the above controller is demonstrated in Sec. V. Some corroborating simulations are provided in Sec. VI. Finally, in Sec. VII we conclude the paper and discuss future directions.
@@ IV-D Platoon safety
As the extensive study of string stability lacks the safety analysis [43, Sec. I], we try to include it in this paper by introducing some surrogate safety measures (SSMs). According to the latest survey [44, Fig. 2], time-based SSMs are the most frequently used SSM for mixed traffic safety assessment, followed by deceleration-based and distance-based SSMs. As there is no generally accepted consensus to classify acceptable and non-acceptable risk level produced by different SSMs [45, Sec. 1], inspired by the simulation in [44, Fig. 4 (c)], we choose three SSMs, i.e., time to collision (TTC), deceleration rate to avoid crash (DRAC) and difference of the space distance and stopping distance (DSS) accordingly.
In detail, TTC and DRAC are defined in [45, Equations (1) and (4)], respectively; DSS is defined in [46, Equation (4)]. These three SSMs are practical and easy to calculate. Denote their thresholds respectively as $TTC^{*},DRAC^{*}$ and 0. It is worth noting that no consensus exists for selecting the SSM thresholds; nonetheless, the outcome of safety analysis relies heavily on threshold selection [44]. Although string stability strictly speaking is not considered as SSM, previous literature has proved that better string stability could bring important safety benefits [47, Sec. 3.3]. As a result, since string stability is considered and solved in this paper (see Sec. V for more details), we choose less conservative thresholds, e.g., $TTC^{*}=1.5s,DRAC^{*}=3.4m/s^{2}$ ($TTC^{*}$ range is typically $1.5s-2s$; $DRAC^{*}$ range is $3m/s^{2}-3.4m/s^{2}$ [44, Sec. III-C]).
When the following three conditions:
$TTC<TTC^{*},DRAC>DRAC^{*},DSS<0$ (17)
are satisfied simultaneously, we mark the situation as unsafe under our observer-based controller. Consequently, the initial condition of the platoon system should not satisfy the unsafe condition (17). Note that the underline relationship between string stability and individual vehicle’s safety is not explicitly understood and needs to be further explored [47, Sec. 4.4], which is one of our future directions.
@@ V String stability
In this section, we first present the string stability condition related to the observer-based controller parameters. Then, a new mechanism is proposed to determine how to design these parameters from the string stability condition using a heuristic searching algorithm. Finally, a time headway minimization algorithm is proposed, which demonstrates that the proposed observer-based controller is capable of further improving the traffic throughput by simulations in Sec. VI-C.
@@ V-B Parameters $k_{1},k_{2},k_{3}$ and $\alpha$ for string stability
Note that from Remark 5, the parameter setting condition (28a) is sufficient but not necessary for string stability condition (23). The advantage, however, is that designing controller parameters becomes much simpler in both calculation and analysis. For example, one method of calculating (23), e.g., in [6], is to get $\left\|H(j\omega)\right\|_{\infty}^{2}\coloneqq\frac{\mathrm{Re}(H_{\text{num}})^{2}+\mathrm{Im}(H_{\text{num}})^{2}}{\mathrm{Re}(H_{\text{den}})^{2}+\mathrm{Im}(H_{\text{den}})^{2}}$ and then, calculate the condition as $(\mathrm{Re}(H_{\text{num}})^{2}+\mathrm{Im}(H_{\text{num}})^{2})-(\mathrm{Re}(H_{\text{den}})^{2}+\mathrm{Im}(H_{\text{den}})^{2})\leq 1$. This method is very calculation-intensive and has more polynomial terms. The details are that $4$ square calculations are needed in the above inequality compared to $2$ in (28a) and the polynomial term related to $\omega^{12}$ will appear in $\mathrm{Re}(H_{\text{den}})^{2}$ while there is no term related to $\omega^{12}$ in (28a) as we can see in (31a). Since the speed of the leader vehicle in [6] is constant, thus one PID controller is designed to stabilize the platoon and parameters for string stability in the above inequality can be calculated as the inequality is simple in [6]. However, in our paper, the leader of varying speed is investigated where observer-based controller is needed to stabilize the platoon; the calculation of $(\mathrm{Re}(H_{\text{num}})^{2}+\mathrm{Im}(H_{\text{num}})^{2})-(\mathrm{Re}(H_{\text{den}})^{2}+\mathrm{Im}(H_{\text{den}})^{2})\leq 1$ will explode and it is very difficult to design parameters accordingly. This is the motivation why a transformed condition (28a) to design parameters was suggested.
@@ V-D Parameter $\alpha$ design simplification and time headway minimization
VI Examples
@@
In this section, the internal stability and the string stability of our proposed observer-based controller are validated. Also, the comparison of the controller ((22) and (14)) in this paper and the following distributed PID controller (36) from [6] is demonstrated:
@@
To demonstrate the string stability, an external disturbance, $u_{0}(t)$ (37) in the following is acted on the leader to have the leader speed profile as shown in Fig. 2 (iii) and (iv):
@@
VI-A Comparison with distributed PID controller (36) in [6]
i Observer error $\xi_{i}$ in (12) between the observer output $\hat{x}_{i}\coloneqq[\hat{p}_{i},\hat{v}_{i},\hat{a}_{i}]^{T}$ and real system unmeasurable states $\tilde{x}_{i}\coloneqq[\tilde{p}_{i},\tilde{v}_{i},\tilde{a}_{i}]^{T}$.
ii Platooning string stability performance comparison for different values of $b$ with $\alpha=1.5,h=0.198s$.
iii The Bode plot of the original string stability condition (23) with different values of $b$ with $\alpha=1.5,h=0.198s$.
iv The relationship between $\left\|H(j\omega)\right\|_{\infty}$ and $b$ when $\alpha=1.5$.
v The relationship between $\left\|H(j\omega)\right\|_{\infty}$ and $\alpha$ when $b=9$.
Figure 3: Heuristic searching algorithm validation for observer-based controller parameters with fixed time headway $h=0.198s$ in Sec. VI-B1.
Based on $r=3$ and $\tau=0.5$, from the complimentary rule (34) in step ii, if we design $\alpha=1.5$, then, the main rule (33a) in step i becomes $30.37.1\leq b<$ 30.3. Based on guidelines, firstly, we design $b=12$. We verify $\{\mathrm{W_{2}},\mathrm{W_{4}},\mathrm{W_{6}}\}\geq 0$ in (32a) as $\mathrm{W_{2}}=5.7773e^{8},\mathrm{W_{4}}=1.0674e^{8},\mathrm{W_{6}}=7.3228e^{6}$ and $\mathrm{W_{6}}+\mathrm{W_{8}}\omega^{2}+\mathrm{W_{10}}(\omega^{2})^{2}\geq 0$ in (32b) for $\omega\in[0,\omega_{0}],\omega_{0}=3.9017$ with $\mathrm{W_{8}}=-1.7811e^{4},\mathrm{W_{10}}=-59.5$. It is worth noting here that it does not mean the transformed string stability condition (31a) is only available for $\omega\in[0,\omega_{0}],$ as we also have $\{\mathrm{W_{2}},\mathrm{W_{4}}\}>0$. However, the calculation for deciding the exact range for $\omega$ with all $\{\mathrm{W_{2}},\mathrm{W_{4}}\}$ included would become very complicated as we can see in (31a). And this is the reason we use step iv to finally and formally confirm our designed feasible parameters $\alpha=1.5$ and $b=12$ as in the Fig. 3 (iii) which means the string stability is guaranteed.
@@ VI-A Comparison with distributed PID controller (36) in [6]
Note that for the presentation convenience, we only plot the speed profiles of vehicles $0,1,3,5,7$; and the distributed PID controller parameters used in this comparing simulation in Table I are from [6]. It is also interesting to find out that in Fig. 2 (iii) with $b=12$, during the initial time period, the speed of vehicle $7$ is negative, which is due to the cause of the peaking effect in Sec. V-C. Based on the main rule (33b), $b=12$ is large. Then, we design $b=9$ and we can see the peaking effect is avoided in Fig. 2 (ii). When we continue decrease the value of $b$ until $b=7.5$ (still satisfies $30.37.1\leq b<$ 30.3), we see no peaking effect but overshoot in the beginning time period in Fig. 2 (i), which also validates our main rule (33b). Note that both $b=9$ and $b=7.5$ satisfy the string stability condition as shown in Fig. 3 (iii).
Furthermore, Fig. 2 (v) and (vi) are used to visualize the string stable performance. Specifically, Fig. 2 (v) shows clearly the internal and string stability of our proposed controller (with $b=9$). Compared with the PID controller in [6] from Fig. 2 (vi), one can see the error magnitude of observer-based controller is much smaller, and its convergence speed is much quicker.
@@
VI-B Analysis of heuristic searching algorithm in Sec. V-C
@@
VI-B1 Condition (32) is satisfied
As we can see, the designed $\alpha$ and $b$ in Sec. VI-A shows (32) is satisfied. In order to further analyze and validate our proposed algorithm, we provide comparison simulations related to the different parameters of the proposed controller, e.g., by keeping $\alpha=1.5$, we give two other choices of $b$ ($b=4$ and $b=35$) which are out of the value range: $30.37.1\leq b<$ 30.3 from the main rule (33a). The Bode plot of different values of parameter $b$ in Fig. 3 (iii) demonstrates that the case of $b=9$ guarantees the string stability while the other two cases do not, which shows the effectiveness of our proposed heuristic searching algorithm (parameter setting mechanism) in Sec. V-C.
More specifically, one can see only Fig. 3 (ii) (b) demonstrates that the platoon is string stable as the error does not amplify from vehicle $1$ to $7$, while the scenarios of $b=4$ in (a) and $b=35$ in (c) are not, which is in accordance with the Bode plot in Fig. 3 (iii).
@@ VI-B1 Condition (32) is satisfied
Figure 4: Feasible region for Sec. VI-B1 and Algorithm 1 for controller parameters with time headway minimization in Sec. VI-C.
It is worth noting that all the errors converge to zero in Fig. 3 (ii) (a-c) and that the error of vehicle $1$ converges to zero first, then that of vehicle $2$, and so on, and finally the error of the last vehicle (vehicle $7$) converges to zero, which verifies Lemma 1 and Theorem 1.
@@
First, assume we set $\alpha=1.5$ without many trials, Fig. 3 (iv) shows that $5.2<b\leq 20.4$ is a decent choice for platooning string stability. One can see this range is different from the range $30.37.1\leq b<$ 30.3 (or even $25.257.1\leq b<$ 25.25 considering the complimentary rule (34) with $b<5/h$), which verifies the above mechanism is neither sufficient nor necessary. However, they do have a quite large overlapped range $7.1\leq b\leq 20.4$ which accounts for $87.5\%$ of $5.2<b\leq 20.4$ and $72.3\%$ of $25.257.1\leq b<$ 25.25, respectively, demonstrating the effectiveness of the above mechanism.
Then, we choose $b=9$ and Fig. 3 (v) shows $0.5<\alpha\leq 3.7$ is acceptable. Therefore, the choice of $(b=9,\alpha=1.5)$ is good to satisfy the string stability condition (23) given the time headway $h$ is fixed. Fig. 4 (i) shows the feasible region of our proposed controller parameter ($\alpha$ and $b$) design mechanism given $h=0.198s$, which also validates the above analysis with $(b=9,\alpha=1.5)$ included.
VI-B2 Condition (32) is not satisfied

### v3

@@ Abstract
In this paper, we study the internal stability and string stability of a vehicle platoon under the constant time headway spacing (CTHS) policy and the multiple-predecessor-following (MPF) vehicle-to-vehicle information flow topology. More specifically, we depart from the conventional Proportional-Integral-Derivative (PID) controller design for such systems and we propose the design of an observer-based controller. For designing our observer-based controller, we first design a distributed observer, with which each follower estimates their position, speed and acceleration error with respect to the leader. The observer is designed by means of constructing an observer matrix whose parameters should be determined. Next, we simplify the design of the matrix of the observer in such a way that the design boils down to choosing a single scalar value; this design further simplifies the structure of the controller, whose simplicity enables the derivation of string stability conditions by means of a frequency response method. Subsequently, the string stability conditions for a given time headway, are transformed to conditions for the controller parameters. We obtain controller parameters that satisfy the stability conditions by designing a novel heuristic search algorithm. Furthermore, we extend the search algorithm by incorporating a bisection-like algorithm, which allows to obtain (within some deviation tolerance) the minimum available value of the time headway. Finally, we provide insights about how to finalize the observer-based controller parameters from above algorithms to avoid the peaking phenomenon. The performance of the proposed observer-based controller, which guarantees internal and string stability, is demonstrated via illustrative examples. Additionally, a comparison with a widely-used PID controller for MPF topology shows that our proposed observer-based controller has better convergence performance. Also, the platoon safety and controller rate convergence are discussed.
@@
Vehicle platooning, varying-speed leader, observer, string stability, safety.
@@ I Introduction
    In order to improve platooning control performance under the MPF topology, a distributed observer-based controller is proposed, which can also ensure the internal and string stability of the platoon system One simulated example in Sec. VII-A demonstrates that this distributed observer-based controller achieves a better platooning control performance compared to the distributed PID controller of [6]. The key point in analyzing string stability is to propose a new calculation mechanism for deriving string stability conditions.
@@
    A bisection-like algorithm is proposed to obtain the minimum (within a tolerance) acceptable value of the time headway that can guarantee string stability. Another simulated example in Sec. VII-C demonstrates that the obtained minimum time headway is smaller than the minimum one from [6] under the same platooning model conditions.
The rest of this paper is organized as follows. Sec. II gives some notations and mathematical preliminaries. Sec. III presents the vehicle model, the spacing policy and the control objective. In Sec. IV, a distributed observer-based controller is proposed with the internal stability analysis. String stability of the above controller is demonstrated in Sec. V. The controller convergence rate is discussed in Sec. VI. Some corroborating simulations are provided in Sec. VII. Finally, in Sec. VIII we conclude the paper and discuss future directions.
@@ IV-D Platoon safety
As the extensive study of string stability lacks the safety analysis [43, Sec. I], we try to include it in this paper by introducing some surrogate safety measures (SSMs). According to the latest survey [44, Fig. 2], time-based SSMs are the most frequently used SSM for mixed traffic safety assessment, followed by deceleration-based and distance-based SSMs. There is no generally accepted consensus to classify acceptable and non-acceptable risk level produced by different SSMs [45, Sec. 1]. Authors in [44] also state that no collective guidance exists to propose the best set of SSMs that can accurately report traffic conflicts. Consequently, inspired by the simulation in [44, Fig. 4 (c)], we choose three SSMs, i.e., time to collision (TTC), deceleration rate to avoid crash (DRAC) and difference of the space distance and stopping distance (DSS) accordingly.
In detail, TTC and DRAC are defined in [45, Equations (1) and (4)], respectively; DSS is defined in [46, Equation (4)]. These three SSMs are practical and easy to calculate. Denote their thresholds respectively as $TTC^{*},DRAC^{*}$ and 0. It is worth noting that no consensus exists for selecting the SSM thresholds; nonetheless, the outcome of safety analysis relies heavily on threshold selection [44]. Although string stability strictly speaking is not considered as SSM, previous literature has proved that better string stability could bring important safety benefits [47, Sec. 3.3]. As a result, since string stability is considered and solved in this paper (see Sec. V for more details), we choose less conservative thresholds, e.g., $TTC^{*}=2s,DRAC^{*}=3.4m/s^{2}$ ($TTC^{*}$ range is typically $1.5s-2s$; $DRAC^{*}$ range is $3m/s^{2}-3.4m/s^{2}$ [44, Sec. III-C]).
When the following three unsafe conditions:
$TTC\leq TTC^{*},DRAC\geq DRAC^{*},DSS\leq 0$ (17)
are satisfied simultaneously, we mark the situation as unsafe under our observer-based controller. Consequently, the initial condition of the platoon system should not satisfy the unsafe condition (17). Note that the underline relationship between string stability and individual vehicle’s safety is not explicitly understood and needs to be further explored [47, Sec. 4.4], which is one of our future research directions.
@@ V String stability
In this section, we first present the string stability condition related to the observer-based controller parameters. Then, a new mechanism is proposed to determine how to design these parameters from the string stability condition using a heuristic searching algorithm. Finally, a time headway minimization algorithm is proposed, which demonstrates that the proposed observer-based controller is capable of further improving the traffic throughput by simulations in Sec. VII-C.
@@ V-B Parameters $k_{1},k_{2},k_{3}$ and $\alpha$ for string stability
Note that from Remark 5, the parameter setting condition (28a) is sufficient but not necessary for string stability condition (23). The advantage, however, is that designing controller parameters becomes much simpler in both calculation and analysis.
@@ V-D Parameter $\alpha$ design simplification and time headway minimization
VI Controller convergence rate discussion
From the controller error dynamics (15) and the observer estimating error dynamics (38), (39) and (40), one can see our controller convergence rate mainly depends on the eigenvalues of $(A-BK)$ with the eigenvalues of $(A-BK-{r_{i}}BL)$ have some influence as well.
As $L=\alpha B^{T}$ in (16), $K=\begin{bmatrix}k_{1}&k_{2}&k_{3}\end{bmatrix}$ in (20) and $k_{1}=b^{3}\tau,k_{2}=3b^{2}\tau,k_{3}=3b\tau-1$ in (30), we have $A-BK=$ $\begin{bmatrix}-b&0&0\\ 0&-b&0\\ 0&0&-b\end{bmatrix},A-BK-{r_{i}}BL=\begin{bmatrix}-b&0&0\\ 0&-b&0\\ 0&0&-b-\frac{r_{i}\alpha}{\tau^{2}}\end{bmatrix}$. We see that parameter $b$ will dominate the convergence rate of our controller. More specifically, the larger the value of $b$ is, the faster the observer-based controller converges.
To sum up, when designing our controller parameters $(\alpha,b)$, in addition to follow the algorithms in Sec. V-C and Sec. V-D, we also take the controller convergence rate into consideration.
VII Examples
@@
Here, the internal stability and the string stability of our observer-based controller are validated. Also, the comparison of our controller ((22) and (14)) in this paper and the following distributed PID controller (36) from [6] is demonstrated:
@@
To demonstrate the string stability, an external disturbance, $u_{0}(t)$ in the following is acted on the leader to have the leader speed profile as shown in Fig. 2 (iii) and (iv):
@@
VII-A Comparison with distributed PID controller (36) in [6]
i Observer error $\xi_{i}$ in (12) between output $\hat{x}_{i}\coloneqq[\hat{p}_{i},\hat{v}_{i},\hat{a}_{i}]^{T}$ and real system unmeasurable states $\tilde{x}_{i}\coloneqq[\tilde{p}_{i},\tilde{v}_{i},\tilde{a}_{i}]^{T}$.
ii Controller convergence rate performance comparison for different values of $b$ with $\alpha=1.5,h=0.198s$.
iii Platooning string stability performance comparison for different values of $b$ with $\alpha=1.5,h=0.198s$.
iv The Bode plot of the original string stability condition (23) with different values of $b$ with $\alpha=1.5,h=0.198s$.
v The relationship between $\left\|H(j\omega)\right\|_{\infty}$ and $b$ when $\alpha=1.5$.
vi The relationship between $\left\|H(j\omega)\right\|_{\infty}$ and $\alpha$ when $b=9$.
Figure 3: Heuristic search algorithm validation for our controller parameters with fixed time headway $h=0.198s$ in Sec. VII-B1.
Based on $r=3$ and $\tau=0.5$, from the complimentary rule (34) in step ii, if we design $\alpha=1.5$, then, the main rule (33a) in step i becomes $30.37.1\leq b<$ 30.3. Based on guidelines, firstly, we design $b=12$. We verify $\{\mathrm{W_{2}},\mathrm{W_{4}},\mathrm{W_{6}}\}\geq 0$ in (32a) as $\mathrm{W_{2}}=5.7773e^{8},\mathrm{W_{4}}=1.0674e^{8},\mathrm{W_{6}}=7.3228e^{6}$ and $\mathrm{W_{6}}+\mathrm{W_{8}}\omega^{2}+\mathrm{W_{10}}(\omega^{2})^{2}\geq 0$ in (32b) for $\omega\in[0,\omega_{0}],\omega_{0}=3.9017$ with $\mathrm{W_{8}}=-1.7811e^{4},\mathrm{W_{10}}=-59.5$. It is worth noting here that it does not mean the transformed string stability condition (31a) is only available for $\omega\in[0,\omega_{0}],$ as we also have $\{\mathrm{W_{2}},\mathrm{W_{4}}\}>0$. However, the calculation for deciding the exact range for $\omega$ with all $\{\mathrm{W_{2}},\mathrm{W_{4}}\}$ included would become very complicated as we can see in (31a). And this is the reason we use step iv to finally and formally confirm our designed feasible parameters $\alpha=1.5$ and $b=12$ as in the Fig. 3 (iv) which means the string stability is guaranteed.
@@ VII-A Comparison with distributed PID controller (36) in [6]
Peaking effect validation. Note that for the presentation convenience, we only plot the speed profiles of vehicles $0,1,3,5,7$. It is also interesting to find out that in Fig. 2 (iii) with $b=12$, during the initial time period, the speed of vehicle $7$ is negative, which is due to the cause of the peaking effect in Sec. V-C. Based on the main rule (33b), $b=12$ is large. Then, we design $b=9$ and we can see the peaking effect is avoided in Fig. 2 (ii). When we continue decrease the value of $b$ until $b=7.5$ (still satisfies $30.37.1\leq b<$ 30.3), we see no peaking effect but overshoot in the beginning time period in Fig. 2 (i), which also validates our main rule (33b). Note that both $b=9$ and $b=7.5$ satisfy the string stability condition as shown in Fig. 3 (iv).
Furthermore, Fig. 2 (v) and (vi) are used to visualize the string stable performance. Specifically, Fig. 2 (v) shows clearly the internal and string stability of our proposed controller (with $b=9$). Compared with the PID controller in [6] from Fig. 2 (vi), one can see the error magnitude of our controller is much smaller, and its convergence speed is much quicker.
@@
Controller convergence rate validation. Fig. 3 (ii) takes vehicle 3 as an example to show the position error convergence rate comparison for different values of $b$ and demonstrates that with a larger value of $b$, our controller converges faster, which validates Sec. VI.
VII-B Analysis of heuristic searching algorithm in Sec. V-C
@@
VII-B1 Condition (32) is satisfied
As we can see, the designed $\alpha$ and $b$ in Sec. VII-A shows (32) is satisfied. In order to further analyze and validate our proposed algorithm, we provide comparison simulations related to the different parameters of the proposed controller, e.g., by keeping $\alpha=1.5$, we give two other choices of $b$ ($b=4$ and $b=35$) which are out of the value range: $30.37.1\leq b<$ 30.3 from the main rule (33a). The Bode plot of different values of parameter $b$ in Fig. 3 (iv) demonstrates that the case of $b=9$ guarantees the string stability while the other two cases do not, which shows the effectiveness of our proposed heuristic searching algorithm (parameter setting mechanism) in Sec. V-C.
More specifically, one can see only Fig. 3 (iii) (b) demonstrates that the platoon is string stable as the error does not amplify from vehicle $1$ to $7$, while the scenarios of $b=4$ in (a) and $b=35$ in (c) are not, which is in accordance with the Bode plot in Fig. 3 (iv).
@@ VII-B1 Condition (32) is satisfied
Figure 4: Feasible region for Sec. VII-B1 and Algorithm 1 for controller parameters with time headway minimization in Sec. VII-C.
It is worth noting that all the errors converge to zero in Fig. 3 (iii) (a-c) and that the error of vehicle $1$ converges to zero first, then that of vehicle $2$, and so on, and finally the error of the last vehicle (vehicle $7$) converges to zero, which verifies Lemma 1 and Theorem 1.
@@
First, assume we set $\alpha=1.5$ without many trials, Fig. 3 (v) shows that $5.2<b\leq 20.4$ is a decent choice for platooning string stability. One can see this range is different from the range $30.37.1\leq b<$ 30.3 (or even $25.257.1\leq b<$ 25.25 considering the complimentary rule (34) with $b<5/h$), which verifies the above mechanism is neither sufficient nor necessary. However, they do have a quite large overlapped range $7.1\leq b\leq 20.4$ which accounts for $87.5\%$ of $5.2<b\leq 20.4$ and $72.3\%$ of $25.257.1\leq b<$ 25.25, respectively, demonstrating the effectiveness of the above mechanism.
Then, we choose $b=9$ and Fig. 3 (vi) shows $0.5<\alpha\leq 3.7$ is acceptable. Therefore, the choice of $(b=9,\alpha=1.5)$ is good to satisfy the string stability condition (23) given the time headway $h$ is fixed. Fig. 4 (i) shows the feasible region of our proposed controller parameter ($\alpha$ and $b$) design mechanism given $h=0.198s$, which also validates the above analysis with $(b=9,\alpha=1.5)$ included.
VII-B2 Condition (32) is not satisfied